I am having a problem with this word problem, I am not good at them and need some help.
Business and finance : The demand eqaurion for a certain type of printer is given by D=200p+35000.
The supply equation is predicted to be S= -P^2+400p-20000
Find the equalibrium price.
Thanks a million

Answer by stanbon(75887)

Business and finance : The demand eqaurion for a certain type of printer is given by D=200p+35000.
The supply equation is predicted to be S= -P^2+400p-20000
Find the equalibrium price.
------------------------
Equilibrium occurs when supply equals demand:
-p^2+400p-20000 = 200p+35000
p^2-200p+15000=0
---------
p = [200+-sqrt(200^2-4*-1*15000)]/(-2)
p = [200+-sqrt(100000)]/(-2)
p = [200+-100sqrt(10)]/(-2)
--------------
Get the positive solution:
p = [200 - 100sqrt(10)]/(-2)
p = $58.11
==========
Cheers,
Stan H.